How long will it take any quantity of iodine 131 to decay to 25% of its initial amount, knowing that it decays according to the function defined by A (t) = A0e^-0.087t, where t is the time in days?
Answer by Boreal(15235)

A (t) = A0e^-0.087t
A(t)/Ao= 0.25, because A(t) is the new level, and that is 25% of the original level.
0.25=e^(-0.087t)
take ln of both sides
-1.38629= -0.087t
t=15.93 days
e^(-0.087*16)=0.2485. This makes sense.